# Quantifying AI Psychology: A Psychometrics Benchmark for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional task-solving capabilities, increasingly adopting roles akin to human-like assistants. The broader integration of LLMs into society has sparked interest in whether they manifest psychological attributes, and whether these attributes are stable—inquiries that could deepen the understanding of their behaviors. Inspired by psychometrics, this paper presents a framework for investigating psychology in LLMs, including psychological dimension identification, assessment dataset curation, and assessment with results validation. Following this framework, we introduce a comprehensive psychometrics benchmark for LLMs that covers six psychological dimensions: personality, values, emotion, theory of mind, motivation, and intelligence. This benchmark includes thirteen datasets featuring diverse scenarios and item types. Our findings indicate that LLMs manifest a broad spectrum of psychological attributes. We also uncover discrepancies between LLMs’ self-reported traits and their behaviors in real-world scenarios. This paper demonstrates a thorough psychometric assessment of LLMs, providing insights into reliable evaluation and potential applications in AI and social sciences.
1 Introduction
@@
The development of large language models (LLMs) has marked a milestone in artificial intelligence (AI) [1, 2, 3]. LLMs demonstrate exceptional performance beyond traditional natural language processing (NLP) tasks [4, 5, 6, 7, 8], with remarkable problem-solving [9, 10, 11] and decision-making abilities [12, 13]. The evolving capabilities of LLMs facilitate their expansion into broader real-world applications [14, 15], directing a significant shift from software tools to general-purpose assistants for humans [16, 17]. It is thus crucial to move beyond merely assessing specific abilities. Inspired by how psychology facilitates the understanding of human behaviors, we investigate psychology in LLMs, aiming to better describe and predict the behaviors of LLMs.
Psychometrics, a systematic evaluation framework, emerges as a promising tool for assessing the psychological attributes of LLMs [18, 19]. It is distinguished by its predictive power and rigorous measurement [20]. Psychometrics evaluates psychological dimensions, termed constructs, which are the hypothesized factors to explain and predict the behaviors of humans [21, 22, 23, 24, 20]. For instance, personality has been shown to predict extensive social outcomes such as career choices and criminal behaviors [25, 26]. Leveraging the predictive power of psychometrics, we intend to identify psychological dimensions and provide insights into the behavioral patterns of LLMs. Additionally, psychometrics emphasizes the importance of evaluation quality by measuring the validity and reliability of the tests [18]. We utilize similar approaches from psychometrics to address concerns about variability in LLMs’ responses and stability of these psychological attributes [27].
@@
As LLMs increasingly fulfill roles as lifelike assistants, there is a growing research interest in quantifying their psychological attributes [28, 29, 30, 31, 32, 33, 34, 35, 36]. Existing evaluations mainly focus on specific dimensions, such as personality [37, 29, 30, 31] or theory of mind [34, 35, 36]. In addition, Miotto et al. [38] provided the initial evidence of psychological assessments for dimensions of personality, values, and demographics in GPT-3. Huang et al. [17] explored psychological portrayals of LLMs, examining dimensions of personality traits, interpersonal relationships, motivational tests, and emotional abilities.
@@
However, there are still two significant challenges that hinder a holistic understanding of LLM psychology. First, existing benchmarks lack diversity and comprehensiveness in both assessment scenarios and item types [38, 17]. Most scenarios only involve self-reported questions (i.e., requiring LLMs to rate themselves), which limits the exploration of their psychological attributes in real-world situations. Additionally, since users primarily interact with LLMs through open-ended questions, it is crucial to understand how these models exhibit their psychological attributes through language generation rather than through closed-form answers. Second, concerns remain about the reliability of the test. These concerns have two aspects: (1) It is unclear whether psychometric tests, designed for humans, are applicable to LLMs. Psychometrics assumes that psychological attributes exist in humans; however, there is a lack of evidence supporting the existence of these attributes in LLMs. For instance, questions arise such as whether LLMs consistently respond to similar situations, whether their preferences for closed-form questions correlate with open-ended responses, and whether their attributes are robust against adversarial attacks; (2) It remains uncertain whether the tests are subject to measurement errors. Besides potential problems caused by position bias [39] and prompt sensitivity [17], our use of LLM-as-a-judge [39] approach for the open-ended responses raises concerns about the reliability of LLM raters. To address these challenges, we present a comprehensive psychometric benchmark for investigating psychology in LLMs, covering six psychological dimensions: personality, values, emotion, theory of mind, motivation, and intelligence.
@@
  • Consistency in responding to similar situations.  LLMs show consistent behavior in tasks that require reasoning, such as theory of mind or emotional intelligence tasks. In contrast, responses to preference-based questions, which do not have definitive answers, display significant variability across different models. Utilizing specific prompts (e.g., role-playing prompts) can improve response consistency toward designated attributes.
@@
  • Closed-form versus open-ended responses.  For instance, a model might score low on extraversion in closed-form assessments yet demonstrate extraverted traits in open-ended responses. This discrepancy is also observed in human responses, where individuals may provide socially desirable answers on rating scales, whereas open-ended questions allow for more nuanced expressions that better reflect complex thoughts [40, 41]. LLMs may simulate responses based on their training data, and open-ended queries might more accurately reveal the model’s underlying generation patterns. This difference can reveal underlying inconsistencies in the model’s learned behavior.
@@
  • Position bias and prompt sensitivity.  The influence of option bias is minimal for models such as GPT-4 and Llama3-70b, whereas it is more apparent in models like ChatGPT and Llama3-8b. Moreover, LLMs exhibit varying degrees of prompt sensitivity during psychometric tests. While most models effectively manage simple substitutions (e.g., noun changes) with minimal impact, logical alterations frequently result in inconsistent outcomes. Additionally, models are particularly susceptible to perturbations in prompts when encountering challenging questions.
@@
  • Reliability of LLM-as-a-judge.  LLM-as-a-judge has been widely used in recent studies [39, 42, 43, 44]. In our study, we utilize two competent LLMs, GPT-4 and Llama3-7b, as raters for evaluating open-ended items. Our examination of their consensus reveals that these LLM raters maintain high consistency across all tests. This consistency demonstrates the potential applicability of this approach in similar evaluation scenarios.
@@
Impact.  Our psychometrics benchmark, situated at the intersection of psychology and AI, carries significant implications for AI, social sciences, and society. We uncover the variability in the behaviors of LLMs across different evaluation scenarios and contexts. These findings deepen our understanding of LLMs’ response patterns, highlighting the need to mitigate biases and undesirable behaviors to develop socially responsible AI [45, 46, 47, 48, 49]. Furthermore, developers can integrate psychological attributes into AI assistants for downstream applications. These lifelike assistants could benefit many sectors, including healthcare, education, and customer service [50, 51]. Our benchmark also serves as a valuable tool for social science research. With the increasing use of LLMs to simulate human responses [52, 53], this benchmark facilitates a more rigorous selection of LLMs for replicating human studies. It further enables interpretable analyses of simulated responses. For the general public, we position LLMs as lifelike assistants that efficiently handle user requests, fostering trust and enhancing the overall user experience.
2 Our Framework of Psychometrics Benchmark
@@
We draw inspiration from psychometrics and employ LLMs as respondents in psychological tests. Though LLMs are trained on extensive datasets that include human opinions and thoughts, it is crucial to acknowledge the fundamental differences between humans and LLMs when conducting psychometrics tests on LLMs. First, the question of whether LLMs possess agency remains debated [54, 55, 56]. The self-reported questionnaires in psychometrics presuppose that respondents possess agency, whereas LLMs’ response reflects a multitude of characters from their training data [57]. Second, LLMs are sensitive to prompt perturbations that humans might find trivial [58, 59]. Acknowledging these differences, we present the framework for the psychometrics benchmark, consisting of three crucial components: psychological dimension identification, assessment dataset curation, and assessment with results validation.
Figure 1: Overview of Our Psychometrics Benchmark for Large Language Models.
2.1 Psychological Dimension Identification
@@
We first need to identify psychological dimensions that could explain and predict the behaviors of LLMs. We adopt a common top-down approach to identify dimensions, i.e., referring to psychology theories and employing an analogy between humans and LLMs [60, 61]. Specifically, we initially draw upon social science and psychology literature as the source of supporting theories for dimension identification. However, this analogy may not always hold due to the fundamental differences between humans and AI models. To bridge this gap, we establish the following guidelines for identifying appropriate psychological dimensions for LLMs:
@@
  • Validity: This guideline suggests that psychological dimensions should be valid constructs to predict behaviors. A counter-example of a valid dimension is astrological signs. Though popular in some cultural contexts for predicting traits, astrological signs lack scientific credibility in psychology and show no consistent impact on human behavior or cognition. In contrast, psychological dimensions that are grounded in scientific theories or empirical evidence possess predictive power that can effectively explain behaviors.
@@
  • Meaningfulness: This guideline asserts that psychological dimensions should be relevant to the capabilities or functions of LLMs that yield meaningful assessment results. For instance, emotional variability is a valid psychological dimension for humans, influencing behaviors in high-stakes environments. However, applying the same concept to LLMs is not meaningful, as emotions in humans arise from biological mechanisms that LLMs do not possess. Conversely, the ability to understand emotion is meaningful for both humans and AI; it enables AI chatbots to comprehend user requests more effectively, thereby enhancing service efficiency.
With the guidelines, we identify six psychological dimensions for investigation: personality, values, emotion, theory of mind, motivation, and intelligence.
2.2 Assessment Dataset Curation
@@
For evaluating these identified psychological dimensions, we curate datasets using three sources as contents: standard psychometrics tests, established datasets, and self-designed scenarios. In total, thirteen datasets (shown in Table 1) are curated with the guidelines detailed in Appx. A. The overview of curated datasets is shown in Table 1.
@@
Table 1: Dataset Overview. “Psych. Test” means Psychometrics test, “Est. Dataset” means Established dataset. ○ means evaluation through the automatic scripts (e.g., keywords matching), ● means the automatic evaluation through LLM-as-a-judge approach using GPT-4 and Llama3-70b as raters.
Dimension Dataset Source Number Item Type Eval
Big Five Inventory [62] Psych. Test 44 Rating-Scale (1~5)
@@
Short Dard Triad [63] Psych. Test 12 Rating-Scale (1~5)
@@
Personality Vignette Test (Big Five) [64] Est. Dataset 5 Open-ended
Cultural Orientation [65] Psych. Test 27 Rating-Scale (1~5)
MoralChoice [66] Est. Dataset 1767 Alternative-Choice
@@
Emotion Understanding [33] Est. Dataset 200 Multiple-Choice
Emotion Emotion Application [33] Est. Dataset 200 Multiple-Choice
False Belief Task [34] Est. Dataset 40 Alternative-Choice
Strange Stories Task [35] Est. Dataset 11 Open-Ended
Theory of Mind Imposing Memory Task [35] Est. Dataset 18 Alternative-Choice
@@
Motivation HoneSet [67] Est. Dataset 987 Open-Ended
These datasets are curated to comprehensively assess each psychological dimension, facilitating an in-depth understanding of LLMs’ behaviors. The construction of each dataset follows the procedure involving content curation, item design, and prompt design:
Content Curation.  The contents of the datasets are sourced from standard psychometric tests or are based on theoretical foundations. These theories not only validate the datasets but also guide the enhancement of dataset diversity. For instance, research on the Theory of Mind (ToM) involves multifaceted tasks encompassing various scenarios and different levels of ToM reasoning. This informs our inclusion of a diverse range of scenarios and reasoning levels in ToM problems.
Item Design.  A key innovation of this benchmark is its capacity to uncover the psychological attributes of LLMs under various evaluation settings, such as self-reported and real-world scenarios. This is achieved by using varied item types to assess a psychological dimension. For instance, to evaluate personality, we incorporate both rating-scale Big Five Inventory and open-ended vignette tests. This approach enables a direct comparison between LLMs’ self-evaluation scores and their narrative responses to real-world scenarios.
2.3 Assessment with Results Validation

### v2

# Evaluating Large Language Models with Psychometrics

@@ Abstract.
Large Language Models (LLMs) have demonstrated exceptional capabilities in solving various tasks, progressively evolving into general-purpose assistants. The increasing integration of LLMs into society has sparked interest in whether they exhibit psychological patterns, and whether these patterns remain consistent across different contexts—questions that could deepen the understanding of their behaviors. Inspired by psychometrics, this paper presents a comprehensive benchmark for quantifying psychological constructs of LLMs, encompassing psychological dimension identification, assessment dataset design, and assessment with results validation. Our work identifies five key psychological constructs—personality, values, emotional intelligence, theory of mind, and self-efficacy —assessed through a suite of 13 datasets featuring diverse scenarios and item types. We uncover significant discrepancies between LLMs’ self-reported traits and their response patterns in real-world scenarios, revealing complexities in their behaviors. Our findings also show that some preference-based tests, originally designed for humans, could not solicit reliable responses from LLMs. This paper offers a thorough psychometric assessment of LLMs, providing insights into reliable evaluation and potential applications in AI and social sciences.
1. Introduction
@@
The development of large language models (LLMs) has marked a milestone in artificial intelligence (AI) (Bommasani et al., 2021; Zhao et al., 2023a). LLMs demonstrate remarkable performance beyond traditional natural language processing (NLP) tasks (Touvron et al., 2023a, b; Qin et al., 2023), with remarkable problem-solving (Yao et al., 2024; Shen et al., 2024) and decision-making abilities (Li et al., 2022a; Shinn et al., 2024). The evolving capabilities of LLMs facilitate their expansion into broader real-world applications (Ma et al., 2023a; Mehandru et al., 2024), directing a significant shift from software tools to general-purpose assistants for humans (Qian et al., 2023; Huang et al., 2024). It is thus crucial to move beyond merely evaluating performance on specific tasks. Inspired by how psychology facilitates the understanding of human behaviors, we investigate psychology in LLMs, aiming to better describe and predict the behaviors of LLMs.
Psychometrics, a scientific discipline that aims to develop and refine quantitative methods to measure latent psychological constructs, emerges as a promising framework for assessing the psychological traits or states of LLMs (Jones and Thissen, 2006; Rust and Golombok, 2014; Huang et al., 2024; Wang et al., 2023a). Latent constructs are the hypothesized factors to explain and predict the behaviors of humans (Embretson and Reise, 2013; Slaney, 2017; Cronbach and Meehl, 1955; Wang et al., 2023a). For instance, personality traits have been shown to predict extensive social outcomes such as career choices and criminal behaviors (Ozer and Benet-Martinez, 2006; Strickhouser et al., 2017). Leveraging the psychometric methodologies, we intend to identify psychological dimensions and provide insights into the behaviors of LLMs. Additionally, psychometrics emphasizes the importance of evaluation quality by measuring the reliability of the responses produced by LLMs (Rust and Golombok, 2014). We extend the psychometric test quality assurance framework to determine whether reliable conclusions can be drawn from LLM responses and to shed light on the sensitivity and variability of LLMs’ behaviors (Xiao et al., 2023).
@@
As LLMs increasingly fulfill roles as general-purpose assistants, there is a growing research interest in quantifying their psychological patterns (Jiang et al., 2024a; Safdari et al., 2023; Huang et al., 2023; Jiang et al., 2023a; Wang et al., 2023b; Sabour et al., 2024; Kosinski, 2023; van Duijn et al., 2023; Wu et al., 2023). Existing evaluations mainly focus on specific dimensions, such as personality (Bodroza et al., 2023; Safdari et al., 2023; Huang et al., 2023; Jiang et al., 2023a) or theory of mind (Kosinski, 2023; van Duijn et al., 2023; Wu et al., 2023). In addition, Miotto et al. (2022) provided the initial efforts of psychological assessments for dimensions of personality, values, and demographics in GPT-3. Huang et al. (2024) explored psychological portrayals of LLMs, examining dimensions of personality traits, interpersonal relationships, motivational tests, and emotional abilities.
However, there are still two challenges that hinder a holistic understanding of LLM psychology:
  • Existing benchmarks lack diversity and comprehensiveness in both assessment scenarios and item types, limiting the analysis of LLM behaviors across various contexts (Miotto et al., 2022; Huang et al., 2024). Most tests only involve self-reported questions (i.e., requiring LLMs to rate themselves), which constrains the exploration of their psychological tendencies in real-world situations. Additionally, since users primarily interact with LLMs through open-ended questions, it is crucial to understand how these models exhibit their psychological patterns through open-ended responses rather than through closed-form answers.
@@
  • Concerns persist regarding the reliability of the tests. These concerns have two aspects: (1) It is unclear whether psychometric tests designed for humans apply to LLMs. Psychometrics assumes the existence of psychological attributes in humans, indicating a certain degree of behavioral consistency. However, there is a lack of evidence supporting the consistency of these psychological patterns in LLMs. For instance, questions arise such as whether LLMs consistently respond to similar situations, whether their preferences for closed-form questions correlate with their responses to open-ended ones, and whether their tendencies remain robust against adversarial attacks; (2) It remains uncertain whether the tests are subject to measurement errors. Besides potential problems caused by position bias (Zheng et al., 2023) and prompt sensitivity (Huang et al., 2024), our use of LLM-as-a-judge (Zheng et al., 2023) approach for the open-ended responses raises concerns about the reliability of LLM raters. To address these challenges, we present a comprehensive psychometric benchmark to investigate psychology in LLMs, which encompasses dimension identification, dataset design, and assessment with results validation. We administer evaluations across five psychological dimensions: personality, values, emotion, theory of mind, and motivation, and discuss how psychometrics can assist in evaluating the intelligence of LLMs.
@@
  • Discrepancies between closed-form and open-ended responses.  LLMs exhibit discrepancies in psychological tendencies when responding to closed-form versus open-ended questions. For example, a model might score low on extraversion in closed-form assessments but display extraversion in open-ended responses. This pattern is also observed in humans, where individuals may provide socially desirable answers on rating scales, while open-ended questions allow for more nuanced expressions that better reflect complex thoughts (Hift, 2014; Baburajan et al., 2022). LLMs may simulate responses based on their training data, and open-ended queries might more accurately reveal the model’s underlying generation patterns. These differences highlight inconsistencies in the model’s learned behavior, suggesting that LLMs lack an internal representation that aligns their self-reported answers with their responses to real-world questions.
  • Consistency in responding to similar situations.  LLMs have consistent performance on tasks that require reasoning, such as theory of mind or emotional intelligence. However, their responses to preference-based questions—those without clear right or wrong answers—vary significantly across different models in similar situations. Some models respond inconsistently to similar situations, making it unreliable to determine the psychological patterns of LLMs based on these responses. Using specific prompts (e.g., role-playing prompts) can improve response consistency toward designated tendencies.
@@
  • Position bias and prompt sensitivity.  The influence of option position bias is almost negligible for models such as GPT-4 and Llama3-70b, whereas it is more prominent in models like ChatGPT and Llama3-8b. Moreover, LLMs exhibit varying degrees of prompt sensitivity in psychometric tests. While most models effectively handle simple substitutions (e.g., noun changes) with minimal impact, logical alterations frequently result in inconsistent outcomes. Additionally, models are particularly susceptible to perturbations in prompts when encountering challenging questions.
@@
  • Reliability of LLM-as-a-judge.  LLM-as-a-judge has been widely used in recent studies (Zheng et al., 2023; Kim et al., 2024). In our study, we employ two capable LLMs, GPT-4 and Llama3-70b, as raters for evaluating open-ended items. Our analysis of their consensus reveals that these LLM raters achieve high agreement across all tests. This agreement demonstrates the potential applicability of this approach in similar evaluation scenarios.
@@
Impact.  Our psychometric benchmark, situated at the intersection of psychology and AI, has significant implications for AI development, social sciences, and society. By revealing variability in LLM behaviors across diverse evaluation scenarios, our findings enhance the understanding of LLM response patterns and emphasize the necessity to mitigate biases for the development of socially responsible AI (Rao et al., 2023; Sun et al., 2024; Gallegos et al., 2024). Additionally, developers can leverage these psychological insights to enhance AI assistants, benefiting sectors such as healthcare, education, and customer service (Kasneci et al., 2023; Yang et al., 2023). For social science research, our benchmark provides a robust tool for selecting appropriate LLMs to simulate human responses (Zhao et al., 2023b; Dillion et al., 2023) and facilitates more interpretable analyses. For the general public, we position LLMs as general-purpose assistants that have the potential to efficiently handle user requests, fostering trust and enhancing the overall user experience.
2. Our Framework of Psychometric Benchmark
@@
Our work links to psychometrics by treating LLMs as respondents in structured evaluations, similar to psychological tests, to analyze their reasoning, consistency, and biases in decision-making tasks. Although LLMs are trained on extensive datasets that encompass human opinions and thoughts, it is essential to recognize the fundamental differences between humans and LLMs when conducting psychometric assessments. First, humans can reflect their genuine feelings and thoughts derived from personal experiences, whereas LLMs lack such mechanisms; LLMs’ responses reflect “a multitude of characters” from their training data (Shanahan et al., 2023). Second, LLMs are highly sensitive to prompt perturbations that humans might find trivial (Lin, 2024; Sclar et al., 2024). Acknowledging these differences, we present our framework for a psychometric benchmark for LLMs, consisting of three crucial components: psychological dimension identification, assessment dataset design, and assessment with results validation, as shown in Fig. 1.
Figure 1. Overview of Our Psychometrics Benchmark for Large Language Models.
2.1. Psychological Dimension Identification
@@
We identify psychological dimensions that could explain and predict the behaviors of LLMs. We adopt a top-down approach to identify dimensions, which involves drawing on psychological theories and analogies between humans and LLMs (Hankin and Abela, 2005; Raykov and Marcoulides, 2011). Specifically, we initially draw upon social science and psychology literature as sources of supporting theories for dimension identification. However, this analogy may not always hold due to the differences between humans and AI models. To bridge this gap, we establish the following guidelines for identifying psychological dimensions for LLMs:
@@
  • Appropriateness: This guideline suggests that psychological dimensions should be appropriate and valid constructs to predict behaviors. One example of an inappropriate dimension is astrological signs. Though popular in some cultural contexts for predicting traits, astrological signs lack scientific credibility in psychology and show no consistent impact on human behavior or cognition. In contrast, psychological dimensions that are grounded in scientific theories or empirical evidence possess predictive power that can effectively explain behaviors.
@@
  • Meaningfulness: This guideline asserts that psychological dimensions should be relevant to the capabilities or functions of LLMs that yield meaningful assessment results. For instance, emotional variability can be a psychological dimension for humans, influencing behaviors in high-stakes environments. However, applying the same concept to LLMs is not meaningful, as emotions in humans arise from biological mechanisms that LLMs do not possess. Conversely, the ability to understand emotion is meaningful for both humans and AI; it enables AI chatbots to comprehend user requests more effectively.
Following these guidelines, we develop datasets to evaluate five psychological dimensions: personality, values, emotion, theory of mind, and motivation. Additionally, we provide a separate discussion on intelligence, an important and well-studied dimension, in Appx. H.
2.2. Assessment Dataset Design
@@
For evaluating these psychological dimensions, we curate datasets using three sources: standard psychometrics tests, established datasets, and self-designed scenarios. In total, 13 datasets (shown in Table 1) are curated with the guidelines detailed in Appx. A. These datasets are curated to comprehensively assess each psychological dimension, facilitating an in-depth understanding of LLMs’ behaviors. The construction of each dataset follows the procedure involving content curation, item design, and prompt design.
Content Curation.  The contents of the datasets are either sourced from standard psychometric tests or based on established theories. These theories not only validate the datasets but also guide the enhancement of dataset diversity. For instance, research on the Theory of Mind (ToM) involves multifaceted tasks encompassing various scenarios and different levels of ToM reasoning. This informs our inclusion of a diverse range of scenarios and reasoning levels in ToM problems.
Item Design.  One innovation of this benchmark is its capacity to uncover the psychological patterns of LLMs under various evaluation settings, such as self-reported and real-world scenarios. This is achieved by using varied item types to assess a psychological dimension. For instance, to evaluate personality, we incorporate both rating-scale Big Five Inventory and open-ended vignette tests. This approach enables a direct comparison between LLMs’ self-evaluation scores and their narrative responses to real-world scenarios.
@@
Table 1. Overview of assessment datasets. “Psych. Test” means Psychometrics test, “Est. Dataset” means Established dataset. ○ indicates evaluation through automatic scripts (e.g., keywords matching), ● indicates automatic evaluation using the LLM-as-a-judge approach, with GPT-4 and Llama3-70b serving as raters.
Dimension Dataset Source # of Items Item Type Eval
Big Five Inventory (John et al., 1999) Psych. Test 44 Rating-Scale (1~5)
@@
Short Dard Triad (Jones and Paulhus, 2014) Psych. Test 12 Rating-Scale (1~5)
@@
Personality Vignette Test (Big Five) (Kwantes et al., 2016) Est. Dataset 5 Open-ended
Cultural Orientation (Hofstede et al., 2010) Psych. Test 27 Rating-Scale (1~5)
MoralChoice (Scherrer et al., 2024) Est. Dataset 1767 Alternative-Choice
@@
Emotion Understanding (Sabour et al., 2024) Est. Dataset 200 Multiple-Choice
Emotion Emotion Application (Sabour et al., 2024) Est. Dataset 200 Multiple-Choice
False Belief Task (Kosinski, 2023) Est. Dataset 40 Alternative-Choice
Strange Stories Task (van Duijn et al., 2023) Est. Dataset 11 Open-Ended
Theory of Mind Imposing Memory Task (van Duijn et al., 2023) Est. Dataset 18 Alternative-Choice
@@
Motivation HoneSet (Gao et al., 2024) Est. Dataset 987 Open-Ended
2.3. Assessment with Results Validation